Grade 11Mechanics

A barrier AB of length 12m is hinged at A. At the lower end, a horizontal spring keep the barrier closed. The height of the water is 6m and the width of the barrier is 5m. Water level is 4m below the hinge A. If minimum elongation of spring to keep the barrier closed is 1m, find the spring constant. [Neglect atmospheric pressure.]

sanskriti srivastava
10 Years agoGrade 11

1 Answer

Vaibhav
8 Years ago
Let point A be hinged pointTorque about A integration ((x-d)gwsxdx)=sgw int. (H^3÷3 -dh^2÷2 +d^3÷6)Putting values; h=12;d=distance of water level from top end;s=density;limit of int.=d to h=72×10^5=k×1 (12)=k=600kN/m